OpenStudy (anonymous):

What is the angular velocity of a 6-foot pendulum that takes 3 seconds to complete an arc of 14.13 feet?

OpenStudy (john_es):

But you can also use the other data using other formulas, \[\Delta\theta=\frac{s}{R}=4.31/1.83=2.36\ rad\] \[\omega=\frac{\Delta\theta}{\Delta t}=2.36/3=0.79\ rad/s\]

OpenStudy (john_es):

All measures in SI.